What is the volume occupied by 28.0 g of oxygen, O2 at STP?
1
Expert's answer
2020-10-26T13:29:12-0400

32 gm of oxygen = 1 mole = 22.4lit (standard result according to gas governing equation)28 gm of oxygen = 28/32 mole So Volume=(28)(22.4)/32 =19.6 lit